1. How much heat is released when 9.22 grams of glucose C6H12O6 reacts according to the following equation?

C6H12O6 + 6O2 --> 6CO2 + 6H2O ∆H = -2803 kJ


2. How much heat is absorbed during photosynthesis when 9.22 grams of glucose C6H12O6 is produced?

6CO2 + 6H2O --> C6H12O6 + 6O2 ∆H = +2803 kJ


3. How much heat is released when 147 grams of NO2(g) is dissolved in excess water?

3NO2 + H2O --> 2HNO3 + NO ∆H = - 138 kJ


4. Calculate the heat released when 74.6 grams of SO2 reacts according to the following equation.

2SO2 + O2 --> 2SO3 ∆H = -99.1 k


1
Expert's answer
2021-06-07T03:12:11-0400

1) -2803 kJ heat released when reacts 180gr/mol glucose, here is reacting 9.22 gr glucose so

∆H= 9.22 gr*(-2803 kJ)/180 gr/mol=-143.57kJ

∆H=-143.57kJ

2) -2803 kJ heat absorbed when produced 180gr/mol glucose, here is producting 9.22 gr glucose so

∆H= 9.22 gr*(+2803 kJ)/180 gr/mol=+143.57kJ

∆H=+143.57kJ

3) ∆H=m(NO2)*∆Hr/3*M(NO2)=147gr*(-138kJ)/3*46g/mol=-147kJ

∆H=-147kJ

4) ∆H=m(SO2)*∆Hr/2*M(SO2)=74.6gr*(-99.1kJ)/2*64gr/mol=-57.75kJ

∆H=-57.75kJ